SHOW WORK

rdered: Robaxin 1 gm PO TID Available: Robaxin 250 mg/tablet Give _______ How many times per day? ______ 7. Ordered: Halcion 0.25 mg PO HS PRN Available: Halcion 0.125 mg/tablet Give _______ How many times per day? ______ 8. Ordered: Digoxin 0.125 mg PO QD Available: Digoxin 0.25 mg/tablet Give:_________ 9. Ordered: Vasotec 7.5 mg PO Q AM Available: Vasotech 5 mg tablet and 10 mg tablet Select the ____ mg tablet and give ____ tablets s
Answer #1

ANSWER.

6) Calculation of number of tablets required per day.

         Ordered dose = 1000 mg (1 gm).

         Available dose = 250 mg/tablet.

   Number of tablets required = Ordered dose/Available dose

                                                = 1000/250

                                               = 4 tablets TID or three times a day.

7) Calculation of number of tablets required.

        Ordered dose = 0.25 mg.

        Available dose = 0.125 mg / tablet.

    Number of tablet required = 0.25/0.125

                                       = 2 tablets hs or one time (at bedtime) a day.

8) Calculation of number of tablets required.

          Ordered dose = 0.125 mg.

         Available dose = 0.25 mg/tablet.

     Number of tablet required = Ordered dose/Available dose

                                               = 0.125/0.25

                                               = 0.5 tablet ( half of a tablet).

9) Calculation of number of tablet required.

          Here

              (select the 5 mg tablet)

         Ordered dose = 7.5 mg

         Available dose = 5 mg/tablet.

Number of tablet required = Ordered dose/Available dose

                                             = 7.5 / 5

                                            = 1 .5 tablet ( one and half of a tablet).